# Successive randomized compression: A randomized algorithm for the compressed MPO–MPS product

### v1

@@ 4.1.1 Basic contract-then-compress
The simplest contract-then-compress method uses the standard SVD-based MPS compression algorithm (see, e.g., [32, Alg. 2]). This scheme is known to produce a near-optimal MPS approximation to $H|\psi\rangle$ [32, Cor. 2.4 and sec. 3], but it is slow, requiring
@@
since it requires converting $H|\psi\rangle$ to canonical form prior to compression. Basic contract-then-compress is typically about $D^{2}\times$ slower than the SRC, in view of 3.3b.
@@ 4.1.2 Randomized contract-then-compress
The paper [41] presents three different randomized MPS compression algorithms; for the purpose of this paper, we consider their randomize-then-orthogonalize method [41, Alg. 3.2]. Using this method, randomized contract-then-compress has a runtime of
@@
This improves on the basic contract-then-compress method, but is still typically a factor $D\times$ slower than SRC. Similar to SRC, randomized contract-then-compress is typically a constant factor less accurate than the basic contract-then-compress method. At present, randomized contract-then-compress is limited to the setting where the output bond dimension $\overline{\chi}$ is provided as an input to the algorithm (i.e., not determined adaptively).
@@ 4.1.3 Variational contract-then-compress
A third possible implementation of contract-then-compress uses variational compression to compress the MPO–MPS product $H|\psi\rangle$. Variational MPS is a cyclic minimization algorithm that attempts to minimize the error $\left\||\phi\rangle-|\eta\rangle\right\|_{\rm F}$ of approximating a provided MPS $|\phi\rangle$ by an MPS $|\eta\rangle$ of smaller dimension; see [42, §4.5.2] and [23, §2.6.2] for details. As with DMRG (and fitting, below), there are one-site and two-site versions of the variational compression method, with the two-site version being somewhat more expensive but adaptive and more robust. With the two-site procedure, variational contract-then-compress costs
@@ 4.2.1 Fitting
The fitting method, proposed by Verstraete and Cirac [28], attempts to solve the optimization problem 4.1 by cyclic minimization, iteratively minimizing over each $\eta^{(i)}$ while holding the other sites. This site-by-site minimization is performed in alternating sweeps left-to-right then right-to-left à la DMRG. See [23, §2.8.2] for details. The fitting method is distinguished from the variational contract-then-compress method because contract-then-compress represents $H|\psi\rangle$ directly as an MPS of bond dimension $D\chi$, which it then compresses; the fitting method avoids ever forming the uncompressed MPO–MPS product $H|\psi\rangle$, leading to a fast algorithm.
The basic Verstraete–Cirac fitting method can suffer from slow convergence and does not admit a natural way to determine the bond dimension of $|\eta\rangle$ adaptively to meet a tolerance. To mitigate these issues, Stoudenmire and White [43] proposed a two-site version of the fitting method which minimizes a pair of sites $\eta^{(i)},\eta^{(i+1)}$ and then splits by SVD, allowing for adaptive determination of the bond dimension. The relation of one-site and two-site fitting is analogous to one-site and two-site DMRG.
On a per-sweep basis, the fitting method is a fast, requiring $\mathcal{O}\left(n\cdot[dD\chi\overline{\chi}^{2}+d^{2}D^{2}\chi^{2}]\right)$ operations. However, as Stoudenmire and White [43] warn, even the more robust two-site fitting algorithm can converge slowly or fail to converge at all, particularly for MPOs that capture long-range interactions. Figure 2 illustrates this behavior for the long range $XY$ Hamiltonian $H$ used in section 5.2. In this example, we compute the MPO–MPS product of $H$ with its ground state, which, again, was obtained using DMRG2 with tolerance $10^{-8}$. Despite performing 10 sweeps, the fitting algorithm fails to converge, or make any progress towards solving the problem.
Figure 2: (Fitting failure mode). Runtime (left) and relative error (right) of the fitting algorithm, SRC, and contract-then-compress on long range XY Hamiltonian given in section 5.2.
@@ 4.3.1 Zip-up
The zip-up method was proposed by Stoudenmire and White [43] as an alternative to the potentially slow convergence of the fitting algorithm. Zip-up begins by left-canonizing the input MPO and MPS at a cost of $\mathcal{O}(n\cdot[dD^{3}+d\chi^{3}])$. Then, it proceeds left-to-right merging sites of the MPO and MPS, truncating as it proceeds. See [23, §2.8.3], [45, §3.2.2] for details.
@@ 4.3.2 Density matrix
The density matrix algorithm is conceptually similar to the randomized method developed in this work. In SRC, we construct an MPS approximation $|\eta\rangle$ to $H|\psi\rangle$ by a sequence of randomized QB approximations $B^{(n)}=H|\psi\rangle\approx\eta^{(n)}\cdot B^{(n-1)}$, $B^{(n-1)}\approx\eta^{(n-1)}\cdot B^{(n-2)}$, etc.; each $B^{(j)}$ is a matrix unfolding of an appropriate tensor network. Instead of randomized QB approximation, the density matrix algorithm constructs each $\eta^{(j)}$ as the leading eigenvectors of the density matrix $\rho^{(j)}\coloneqq B^{(j)}(B^{(j)})^{\dagger}$.
@@
The main disadvantage of the density matrix algorithm is speed, being the second slowest method in Figure 1. SRC is nearly as accurate as the density matrix algorithm and often much faster. Another issue with the density matrix algorithm is numerical precision. Forming density matrix $B^{(j)}(B^{(j)})^{\dagger}$ effectively squares the matrix $B^{(j)}$, potentially halving the effective numerical precision. On certain examples, the relative error of the density matrix can stagnate near $10^{-8}$ in double precision ($10^{-4}$ in single), where other methods achieve full accuracy ($\approx 10^{-16}$ in double, $10^{-8}$ in single). The density matrix algorithm faces these precision issues only on some examples, achieving full precision on others; we do not have an explanation for this inconsistency.
@@ 4.4 Comparison
In this section, we compare the SRC method to these alternatives and provide guidance about which algorithm to use.
Table 1: Simplified operation counts for methods to apply an MPO (bond dimension $D$) to an MPS (bond dimension $\chi$). For this table, we assume $d=\mathcal{O}(1)$ and $D\leq\chi\leq\overline{\chi}$; see Table 2 for a complete operation count that does not make these assumptions.
@@ 4.4.1 Complexity
We first compare the methods based on their operation counts, which are shown in Table 1. This table displays simplified operation counts where the physical dimension $d$ is treated as a constant and the bond dimensions are ordered as $D\leq\chi\leq\overline{\chi}$; see Table 2 for more detailed operation counts in terms of all five parameters $n$, $d$, $\chi$, $D$, and $\overline{\chi}$.
In terms of pure operation count, we see that the randomized and zip-up methods are fastest followed by density matrix and randomized contract-then-compress, all of which are faster than basic contract-then-compress. These operation counts are consistent with the runtime experiments provided in paper. The per-iteration cost of the fitting method is also comparable to the randomized and zip-up methods, but—as Figure 2 shows—the fitting method can take many iterations to converge or fail to converge outright for difficult examples. Finally, we notice that the operation counts of all methods are comparable when $\overline{\chi}\approx D\chi$; therefore, speedups are only possible for MPO–MPS multiplication when a significant amount of compression is used $\overline{\chi}\ll D\chi$.
Figure 3: (Fixed bond dimension). Runtime (left) and relative error (right) of several MPO–MPS multiplication algorithms for a synthetic problem as a function of the output bond dimension $\overline{\chi}$, as described in section 4.4.2.
@@ 4.4.2 Comparison experiment 1: Fixed output bond dimension
First, we compare the methods when the output bond dimension $\overline{\chi}$ is fixed by the user. We use the same problem setup as section 1, where we generate random a random MPS and MPO with bond dimensions $D=\chi=50$ across $n=100$ sites and uniformly random entries $[\alpha,1]$ with $\alpha=-0.5$. As in the introduction, we use complex data types and use the contract-then-compress method with a machine-precision tolerance to compute a baseline.
A comparison of existing MPO–MPS methods and SRC is presented in Figure 3 for a range of requested bond dimensions $\overline{\chi}$. We implement SRC both as-is and with the oversampling strategy presented in section 3.4. For this problem, we see that SRC is the fastest method for all requested output bond dimensions $\overline{\chi}$ and, when run with oversampling, achieves comparable accuracy to contract-then-compress at a faster runtime than other accurate methods.
@@ 4.4.3 Comparison experiment 2: Adaptive determination of bond dimension
Next, we compare the the methods when the bond dimension is determined adaptively to meet a tolerance; see appendix C for how to implement SRC in this setting. The runtime behavior of the algorithms in this setting is somewhat changed, as the adaptive determination of bond dimension places a different level of overhead on each algorithm.
Figure 4 shows the results. For both the density matrix and contract-then-compress methods, the runtime behavior remains largely unaffected by the requested tolerance. SRC continues to be the fastest method in this setting, except at large values of the tolerance parameter where the zip-up method becomes fastest. This speed of zip-up comes with a cost, with the zip-up method resulting in meaningfully larger bond dimensions than other algorithms. Therefore, even for high values of the tolerance, SRC (with a final round) may be preferable to obtain an output of the smallest possible bond dimension, with the zip-up method being preferable for fast, approximate computations.
@@ 5 Application: Unitary time evolution
To demonstrate the effectiveness of the randomized contraction algorithm, we compare the performance of several compressed MPO–MPS multiplication algorithms for computing the unitary time evolution $|\psi(t)\rangle=\mathrm{e}^{-\mathrm{i}tH}|\psi(0)\rangle$ of an initial state $|\psi(0)\rangle$ under the influence of a Hamiltonian $H$. Section 5.1 describes the TDVP-GSE algorithm, a leading algorithm for time evolution that requires compressed MPO–MPS products, and section 5.2 provides a comparison of different MPO–MPS algorithms for use with this algorithm.
@@ 5.1 The time-dependent variational principle with global subspace expansion
The time-dependent variational principle (TDVP) algorithm [46] and its variants are currently among the best-performing methods for simulating time-evolution. In its basic form (“TDVP1”), TDVP simulates time evolution by updating individual site of the MPS in sequential left-to-right, right-to-left sweeps. This basic one-site procedure can have issues (similar to one-site DMRG), and even two-site versions of TDVP (“TDVP2”, [47, App. 5]) can struggle on difficult examples. To address this issues, Yang and White [27] proposed an enhancement to TDVP by enriching the state at each time step with a set of global Krylov vectors. Given a current state $|\psi(t)\rangle$ and timestep $\delta$, a single step of Yang and White’s method proceeds as follows:
@@ 5.2 Comparison of MPO–MPS methods for GSE-TDVP1
To simulate this system, we represent $H$ as an MPO using the linear-combination-of-exponentials method of [48, App] with tolerance $10^{-8}$, and we compute the ground state $|\psi_{\rm min}\rangle$ using DMRG2 with tolerance $10^{-8}$. The bond dimensions of $H$ and $|\psi(0)\rangle$ are $D=26$ and $\chi=32$. Time evolution is performed with GSE-TDVP1 with a time step of $\delta=0.2/J$ to a final time $tJ=5$, and all MPO–MPS products $(\mathrm{I}-\delta H)|\phi\rangle$ are truncated to have the same bond dimension as $|\phi\rangle$. Our purpose in presenting this experiment is to compare several MPO–MPS methods for use in GSE-TDVP1, not to claim that the GSE-TDVP1 method is necessary to resolve this particular physical system. (Indeed, the qualitative features of this model can be resolved even using TDVP1.)
Figure 5 shows the results. The left panel shows that the magnetization difference 5.1 exhibits a light cone with power law leakage, in concurrence with the results of previous studies. The middle and right panels show the runtime with various MPO–MPS methods. (The fitting method yielded very poor accuracy and resulted in a rapid explosion in the bond dimension during the basis expansion step, so we omit it.) The SRC method is definitely the fastest, achieving speedups of up to 181$\times$ over contract-then-compress, 45$\times$ over density matrix, and 3.2$\times$ over zip-up for the Krylov vector construction.
Figure 5: (TDVP unitary time evolution). Left: Magnetization difference 5.1 at each position $1\leq i\leq 101$ and time $0\leq tJ\leq 5$ computed by GSE-TDVP1 with randomized MPO–MPS multiplication, showing a light cone with power law leakage. Middle and right: Total runtime (middle) and runtime spent during Krylov vector construction (right) for different MPO–MPS methods per timestep.
@@ Acknowledgements
Appendix A Proof of Theorem 2
@@
Since $A$ is rank-$p$, $A$ contains $p$ linearly independent columns $A(:,i_{1}),\ldots,A(:,i_{p})$. The matrix $\Omega=\begin{bmatrix}e^{(i_{1})}&\cdots&e^{(i_{p})}\end{bmatrix}$ can be realized as a Khatri–Rao product $\Omega=\Omega^{(1)}\odot\cdots\odot\Omega^{(n)}$, where $e^{(i)}$ denotes the standard basis vector with entries $e^{(i)}_{j}=\delta_{ij}$. Thus, $p(\Omega^{(1)},\ldots,\Omega^{(n)})=\det(A(:,\{i_{1},\ldots,i_{p}\})^{\dagger}A(:,\{i_{1},\ldots,i_{p}\}))>0$, completing the proof. $\blacksquare$
Appendix B Proof of Theorem 3
At first glance, it might seem that Theorem 3 is a trivial consequence of Theorem 2, which guarantees rank-$r$ QB decomposition is exact with probability one when applied to a matrix of rank at most $r$. The weakness in this argument is that SRC uses a common set of matrices $\Omega^{(1)},\ldots,\Omega^{(n-1)}$ throughout the algorithm, and the matrix which we applied QB decomposition at each step depends on previous steps of the algorithm. The dependencies between the steps of the algorithm could be a potential source of trouble.

### v2

@@ 4.1.1 Basic contract-then-compress
The simplest contract-then-compress method uses the standard SVD-based MPS compression algorithm (see, e.g., [34, Alg. 2]). This scheme is known to produce a near-optimal MPS approximation to $H|\psi\rangle$ [34, Cor. 2.4 and sec. 3], but it is slow, requiring
@@
since it requires converting $H|\psi\rangle$ to canonical form prior to compression. Basic contract-then-compress is typically about $D^{2}$ times slower than the SRC, in view of ˜3.3b.
@@ 4.1.2 Randomized contract-then-compress
The paper [51] presents three different randomized MPS compression algorithms; for the purpose of this paper, we consider their randomize-then-orthogonalize method [51, Alg. 3.2]. Using this method, randomized contract-then-compress has a runtime of
@@
This improves on the basic contract-then-compress method, but is still typically a factor $D$ times slower than SRC. Similar to SRC, randomized contract-then-compress is typically a constant factor less accurate than the basic contract-then-compress method. At present, randomized contract-then-compress is limited to the setting where the output bond dimension $\overline{\chi}$ is provided as an input to the algorithm (i.e., not determined adaptively).
@@ 4.1.3 Variational contract-then-compress
A third possible implementation of contract-then-compress uses variational compression to compress the MPO–MPS product $H|\psi\rangle$. Variational MPS is a cyclic minimization algorithm that attempts to minimize the error $\left\||\phi\rangle-|\eta\rangle\right\|_{\rm F}$ of approximating a provided MPS $|\phi\rangle$ by an MPS $|\eta\rangle$ of smaller dimension; see [57, §4.5.2] and [23, §2.6.2] for details. As with DMRG (and fitting, below), there are one-site and two-site versions of the variational compression method, with the two-site version being somewhat more expensive but adaptive and more robust. With the two-site procedure, variational contract-then-compress costs
@@ 4.2.1 Fitting
The fitting method, proposed by Verstraete and Cirac [28], attempts to solve the optimization problem ˜4.1 by cyclic minimization, iteratively minimizing over each $\eta^{(i)}$ while holding the other sites. This site-by-site minimization is performed in alternating sweeps left-to-right then right-to-left, à la DMRG. See [23, §2.8.2] for details. The fitting method is distinguished from the variational contract-then-compress method because contract-then-compress represents $H|\psi\rangle$ directly as an MPS of bond dimension $D\chi$, which it then compresses; the fitting method avoids ever forming the uncompressed MPO–MPS product $H|\psi\rangle$, leading to a fast algorithm.
The basic Verstraete–Cirac fitting method can suffer from slow convergence and does not admit a natural way to determine the bond dimension of $|\eta\rangle$ adaptively to meet a tolerance. To mitigate these issues, Stoudenmire and White [58] proposed a two-site version of the fitting method which minimizes a pair of sites $\eta^{(i)},\eta^{(i+1)}$ and then splits by SVD, allowing for adaptive determination of the bond dimension. The relation of one-site and two-site fitting is analogous to one-site and two-site DMRG. All experiments in this paper use the two-site fitting algorithm.
On a per-sweep basis, the fitting method is a fast, requiring $\mathcal{O}\left(n\cdot[dD\chi\overline{\chi}^{2}+d^{2}D^{2}\chi^{2}]\right)$ operations. However, as Stoudenmire and White [58] warn, even the more robust two-site fitting algorithm can converge slowly or fail to converge at all, particularly for MPOs that capture long-range interactions. Figure˜2 illustrates this behavior for the long range $XY$ Hamiltonian $H$ used in section˜5.2. In this example, we compute the MPO–MPS product of $H$ with its ground state $|\psi_{0}\rangle$, which, again, was obtained using DMRG2 with tolerance $10^{-8}$. The fitting algorithm is supplied with the ground state $|\psi_{0}\rangle$ as the initial guess, which agrees with the output $H|\psi_{0}\rangle$ up to scaling by the energy $E_{0}\neq 0$ and numerical errors. Despite performing 10 sweeps, the fitting algorithm fails to converge, or make any progress towards solving the problem. We observe the same behavior when we initialize the fitting algorithm with a random MPS with Gaussian random entries.
Figure 2: (Fitting failure mode). Runtime (left) and relative error (right) of the fitting algorithm, SRC, and contract-then-compress on the MPO–MPS product between the long range XY Hamiltonian given in section˜5.2 and its groundstate.
Similar to the fitting algorithm, the DMRG algorithm can also suffer from convergence to local minima. Researchers have suggested adding random noise [59, 60] or using randomized low-rank approximation algorithms [48] as strategies to avoid this behavior, and these strategies can also be applied to the fitting algorithm. We have not evaluated these more sophisticated versions of the fitting algorithm in this paper.
@@ 4.3.1 Zip-up
The zip-up method was proposed by Stoudenmire and White [58] as an alternative to the potentially slow convergence of the fitting algorithm. Zip-up begins by left-canonizing the input MPO and MPS at a cost of $\mathcal{O}(n\cdot[dD^{3}+d\chi^{3}])$. Then, it proceeds left-to-right merging sites of the MPO and MPS, truncating as it proceeds. See [23, §2.8.3], [62, §3.2.2] for details.
@@ 4.3.2 Density matrix
The density matrix algorithm [63, 64] is conceptually similar to the randomized method developed in this work. In SRC, we construct an MPS approximation $|\eta\rangle$ to $H|\psi\rangle$ by a sequence of randomized QB approximations $B^{(n)}=H|\psi\rangle\approx\eta^{(n)}\cdot B^{(n-1)}$, $B^{(n-1)}\approx\eta^{(n-1)}\cdot B^{(n-2)}$, etc.; each $B^{(j)}$ is a matrix unfolding of an appropriate tensor network. Instead of randomized QB approximation, the density matrix algorithm constructs each $\eta^{(j)}$ as the leading eigenvectors of the density matrix $\rho^{(j)}\coloneqq B^{(j)}(B^{(j)})^{\dagger}$.
@@
The main disadvantage of the density matrix algorithm is speed, being the second slowest method in Figure˜1. SRC is nearly as accurate as the density matrix algorithm and often much faster. Another issue with the density matrix algorithm is numerical precision. Forming density matrix $B^{(j)}(B^{(j)})^{\dagger}$ effectively squares the matrix $B^{(j)}$, potentially halving the effective numerical precision. On certain examples, the relative error of the density matrix can stagnate near $10^{-8}$ in double precision ($10^{-4}$ in single), where other methods achieve full accuracy ($\approx 10^{-16}$ in double, $10^{-8}$ in single). The density matrix algorithm faces these precision issues only on some examples, achieving full precision on others; we do not have an explanation for this inconsistency.
@@ 4.4 Comparison
In this section, we compare the SRC method to these alternatives and provide guidance about which algorithm to use. In all experiments, we use real standard Gaussian entries in random Khatri–Rao products for the SRC algorithm.
Table 1: Simplified operation counts for methods to apply an MPO (bond dimension $D$) to an MPS (bond dimension $\chi$). For this table, we assume $d=\mathcal{O}(1)$ and $D\leq\chi\leq\overline{\chi}\leq D\chi$; see Table˜2 for a complete operation count that does not make these assumptions.
@@ 4.4.1 Complexity
We first compare the methods based on their operation counts, which are shown in Table˜1. This table displays simplified operation counts where the physical dimension $d$ is treated as a constant and the bond dimensions are ordered as $D\leq\chi\leq\overline{\chi}$; see Table˜2 for more detailed operation counts in terms of all five parameters $n$, $d$, $\chi$, $D$, and $\overline{\chi}$.
In terms of pure operation count, we see that the randomized and zip-up methods are fastest followed by density matrix and randomized contract-then-compress, all of which are faster than basic contract-then-compress. These operation counts are consistent with the runtime experiments provided in paper. The per-iteration cost of the fitting method is also comparable to the randomized and zip-up methods, but—as Figure˜2 shows—the fitting method can take many iterations to converge or fail to converge outright for difficult examples. Finally, we notice that the operation counts of all methods are comparable when $\overline{\chi}\approx D\chi$; therefore, speedups are only possible for MPO–MPS multiplication when a significant amount of compression is used $\overline{\chi}\ll D\chi$.
Figure 3: (Fixed bond dimension). Runtime (left) and relative error (right) of several MPO–MPS multiplication algorithms for a synthetic problem as a function of the output bond dimension $\overline{\chi}$, as described in section˜4.4.2.
@@ 4.4.2 Comparison experiment 1: Fixed output bond dimension
First, we compare the methods when the output bond dimension $\overline{\chi}$ is fixed by the user. We use the same problem setup as section˜1, where we generate random a random MPS and MPO with bond dimensions $D=\chi=50$ across $n=100$ sites and uniformly random entries $\alpha,1]$ with $\alpha=-0.5$. As in the introduction, we use complex data types and apply the contract-then-compress method with a machine-precision tolerance to compute a baseline.
A comparison of existing MPO–MPS methods and SRC is presented in Figure˜3 for a range of requested bond dimensions $\overline{\chi}$. We implement SRC both as-is and with the oversampling strategy presented in section˜3.4. For this problem, we see that SRC is the fastest method for all requested output bond dimensions $\overline{\chi}$ and, when run with oversampling, achieves comparable accuracy to contract-then-compress at a faster runtime than other accurate methods.
@@ 4.4.3 Comparison experiment 2: Adaptive determination of bond dimension
Next, we compare the the methods when the bond dimension is determined adaptively to meet a tolerance; see appendix˜C for how to implement SRC in this setting. The runtime behavior of the algorithms in this setting is somewhat changed, as the adaptive determination of bond dimension places a different level of overhead on each algorithm.
Figure˜4 shows the results. For both the density matrix and contract-then-compress methods, the runtime behavior remains largely unaffected by the requested tolerance. SRC continues to be the fastest method in this setting, except at large values of the tolerance parameter where the zip-up method becomes fastest. This speed of zip-up comes with a cost, with the zip-up method resulting in meaningfully larger bond dimensions than other algorithms. Therefore, even for high values of the tolerance, SRC (with a final round) may be preferable to obtain an output of the smallest possible bond dimension, with the zip-up method being preferable for fast, approximate computations.
@@ 5 Application: Unitary time evolution
To demonstrate the effectiveness of the randomized contraction algorithm, we compare the performance of several compressed MPO–MPS multiplication algorithms for computing the unitary time evolution $|\psi(t)\rangle=\mathrm{e}^{-\mathrm{i}tH}|\psi(0)\rangle$ of an initial state $|\psi(0)\rangle$ under the influence of a Hamiltonian $H$. Section˜5.1 describes the TDVP-GSE algorithm, a leading algorithm for time evolution that requires compressed MPO–MPS products, and section˜5.2 provides a comparison of different MPO–MPS algorithms for use with this algorithm.
@@ 5.1 The time-dependent variational principle with global subspace expansion
The time-dependent variational principle (TDVP) algorithm [65] and its variants are currently among the best-performing methods for simulating time-evolution. In its basic form (“TDVP1”), TDVP simulates time evolution by updating individual site of the MPS in sequential left-to-right, right-to-left sweeps. This basic one-site procedure can have issues (similar to one-site DMRG), and even two-site versions of TDVP (“TDVP2”, [66, App. 5]) can struggle on difficult examples. To address this issues, Yang and White [27] proposed an enhancement to TDVP by enriching the state at each time step with a set of global Krylov vectors. Given a current state $|\psi(t)\rangle$ and timestep $\delta$, a single step of Yang and White’s method proceeds as follows:
@@ 5.2 Comparison of MPO–MPS methods for GSE-TDVP1
To simulate this system, we represent $H$ as an MPO using the linear-combination-of-exponentials method of [67, App] with tolerance $10^{-8}$, and we compute the ground state $|\psi_{\rm min}\rangle$ using DMRG2 with tolerance $10^{-8}$. The bond dimensions of $H$ and $|\psi(0)\rangle$ are $D=26$ and $\chi=32$. Time evolution is performed with GSE-TDVP1 with a time step of $\delta=0.2/J$ to a final time $tJ=5$, and all MPO–MPS products $(\mathrm{I}-\delta H)|\phi\rangle$ are truncated to have the same bond dimension as $|\phi\rangle$. Our purpose in presenting this experiment is to compare several MPO–MPS methods for use in GSE-TDVP1, not to claim that the GSE-TDVP1 method is necessary to resolve this particular physical system. (Indeed, the qualitative features of this model can be resolved even using TDVP1.)
[…]
Figure 5: (TDVP unitary time evolution). Left: Magnetization difference ˜5.1 at each position $1\leq i\leq 101$ and time $0\leq tJ\leq 5$ computed by GSE-TDVP1 with randomized MPO–MPS multiplication, showing a light cone with power law leakage. Middle and right: Total runtime (middle) and runtime spent during Krylov vector construction (right) for different MPO–MPS methods per timestep.
@@ Acknowledgements
Appendix A Proof of Theorem˜2
@@
Since $A$ is rank-$p$, AA contains $p$ linearly independent columns $A(:,i_{1}),\ldots,A(:,i_{p})$. The matrix $\Omega=\begin{bmatrix}e^{(i_{1})}&\cdots&e^{(i_{p})}\end{bmatrix}$ can be realized as a Khatri–Rao product $\Omega=\Omega^{(1)}\odot\cdots\odot\Omega^{(n)}$, where $e^{(i)}$ denotes the standard basis vector with entries $e^{(i)}_{j}=\delta_{ij}$. Thus, $p(\Omega^{(1)},\ldots,\Omega^{(n)})=\det(A(:,\{i_{1},\ldots,i_{p}\})^{\dagger}A(:,\{i_{1},\ldots,i_{p}\}))>0$, completing the proof. $\blacksquare$
Appendix B Proof of Theorem˜3
At first glance, it might seem that Theorem˜3 is a trivial consequence of Theorem˜2, which guarantees rank-$r$ QB decomposition is exact with probability one when applied to a matrix of rank at most $r$. The weakness in this argument is that SRC uses a common set of matrices $\Omega^{(1)},\ldots,\Omega^{(n-1)}$ throughout the algorithm, and the matrix which we applied QB decomposition at each step depends on previous steps of the algorithm. The dependencies between the steps of the algorithm could be a potential source of trouble.